OpenStudy (anonymous):

can 4x^2-3y^2 factorize? pls help im stuck

OpenStudy (anonymous):

It can, but you have root 3 y in there.

OpenStudy (anonymous):

? sorry I don't understand what do you mena by root 3y

OpenStudy (unklerhaukus):

\[a^2-b^2=(a+b)(a-b)\] \[4x^2-3y^2=(2x)^2-(\sqrt3 y)^2\]

OpenStudy (anonymous):

:( i'm not getting it, sigh so frustrating.

OpenStudy (christos):

It is basically a simple "identity" generally the rule you are using here to find this is a^2 - b^2 = (a - b) (b - a) Do you get this so far?

OpenStudy (anonymous):

@Christos yeah i learned that bit so far but I cannot find the right combination. we were taught that to factorize something it looks like (a+b)(a-b) but i don't get how it becomes (2x)^2 - (sqroot of 3y)^2 ?

OpenStudy (anonymous):

nvm i get how it becomes that... but, that form wasn't taught to us. we were always taught the (a+b)(a-b) form which we then use to divide...... does (2x)^2 - (sqroot3)^2 change into a (a+b)(a-b) form?

OpenStudy (anonymous):

You've formatted it as a^2-b^2, so yes.

OpenStudy (anonymous):

ok so basically it's just the sq root of first term - sq root of seond term x sq root of first term + sq root of second term?

OpenStudy (unklerhaukus):

yeah